# MOSAIC: Generating Consistent, Privacy-Preserving Scenes from Multiple Depth Views in Multi-Room Environments

### v2

@@ 4.4 Pixel Space Refinement Loss
While our projection loss effectively ensures consistency in latent space, the non-linear nature of the VAE [17] decoder means that consistency in latent space does not necessarily translate to pixel-space consistency. This discrepancy is particularly pronounced when projection transformations are far from orthogonal. We address this issue through a pixel space refinement process during the final denoising stages. For each view $i$, we decode its predicted latent into pixel space: $\hat{x}^{[i]}=g(\hat{z}_{0}^{[i]})$. It is infeasible to calculate $L_{Proj}$ in pixel space, because back-propagation through multi-channel VAE decoder poses computation difficulties. We propose to warp these decoded images between views and compute representations for each view point by depth-weighted fusing the overlapped views:
@@
where $\alpha$ control the selectivity of weighting and $\pi_{ij}$ is the image warping from $j$ view to $i$ view. These optimal pixel-space representation are re-encoded to latent space: ${z}_{0}^{[i]*}=\beta\cdot f({x}^{[i]*})$. The Pixel Space Refinement Loss is defined as:
@@ 4.5 Properties of MOSAIC
Table 1: Quantitative evaluation with Kernel Inception Distance (KID), CLIPIQA+, CLIPScore (CS) and CLIPConsistency(CC).
Method KID $\downarrow$ CS $\uparrow$ CIQA $\uparrow$ CC $\uparrow$
Warp-Inpaint [21] 0.04646 0.6849 0.6517 29.93
MVDiffusion [34] 0.03640 0.7016 0.6025 29.41
SceneCraft [39] 0.07697 0.7011 0.4531 27.68
Text2Room [13] 0.08594 0.6618 0.4388 27.34
@@
5 Experimental Setup
Navigation Agent Autonomous Data Collection. To obtain testing data, we deployed an indoor navigation robot to explore various indoor scenes, collecting depth information via onboard sensors while recording precise camera poses along its trajectory. This approach ensures aligned trajectory-based inputs with baseline methods [34, 39, 13, 21]. To evaluate generalizability, we gathered data across 16 HM3D [28] and 5 MP3D [6] scenes. For trajectory sample details, please refer to Appendix. To generate captions for these depth images, we visualized the collected depth data and utilized a VLM: ChatGPT-4o [5] to provide contextually relevant scene descriptions. This results in 2011 depth data, camera pose, caption pairs in total.
Baselines. We compare our approach against state-of-the-art methods of multiview image generation with geometric priors, including MVDiffusion [34], Warp-Inpaint [21], and SceneCraft [39]. Although SceneCraft shares different condition settings, where its input is the semantic map and the depth map of the bounding box, we create the layout of the bounding box according to our dataset for fair comparison. We evaluate semantic quality against Text2Room [13], a specialized method for indoor scene generation.
Evaluation metrics. The goal of multiview image generation with depth prior is to produce images that maintain consistency in overlapping regions, while preserving image quality, style coherence, and accurate geometric alignment. To evaluate image consistency, we follow MVDiffusion [34] by calculating Warped PSNR [14] and Warped Ratio. We assess image quality using Kernel Inception Distance (KID) [3] and CLIPIQA+ [36], and measure image-text alignment with CLIPScore [10] and CLIPConsistency [27].
Figure 3: Qualitative comparison with multi-view generation baselines. More results are available in the supplementary material.
@@ 6.1 Evaluation Against Baselines
As shown in Fig. 3, MOSAIC produces more coherent and photorealistic results than existing approaches using identical depth inputs. MVDiffusion [34] lacks cross-view consistency and cannot generalize to stylistic prompts like “in Van Gogh style.” The warp-and-inpainting approach [21] generates acceptable initial frames but suffers from progressive style drift and accumulating errors during sequential generation. SceneCraft [39] fails to capture complex environment layouts and produces blurry artifacts due to its NeRF-based methodology. Text2Room [13], built on the warp-and-inpainting pipeline, exhibits noticeable patchwork-like seams from inpainting inconsistencies as well as the depth misalignment. In contrast, MOSAIC maintains consistency across viewpoints while preserving fidelity to both depth information and style prompts, effectively eliminating boundary artifacts through our depth-weighted projection mechanism.
Quantitative Comparisons.
Table 2: Cross-view consistency analysis.
@@
MVDiffusion 13.58 0.53
Warp-Inpaint 22.00 0.86
Ours Pixel 25.30 0.99
Tab. 1 presents a quantitative comparison of our proposed method against several state-of-the-art approaches, including Warp-and-Inpainting [21], MVDiffusion [34], SceneCraft [39], and Text2Room [13], on the multi-view consistent image generation task. Our method achieves the lowest KID score ($\downarrow$) of 0.0391 and highest CS ($\uparrow$) of 0.7166, CIQA of 0.6526, and CC of 30.85, outperforming other methods by a significant margin in all image quality metrics. In addition to the quality improvement, Tab. 2 demonstrates our exceptional cross-view consistency with a Warped PSNR of 25.30 and Warped Ratio of 0.99, approaching ground truth performance (25.45, 1.00) and substantially outperforming competing methods. Although MVDiffusion achieves a competitive KID score of 0.0640, its cross-view consistency is highly limited with a Warped PSNR of only 13.58. Our approach achieves outstanding performance in both image quality and cross-view consistency compared with state-of-the-art methods.
@@ 6.2 Ablation Study
Figure 4: Qualitative ablation results. Red: artifacts; orange: pixel space texture misaligned; Green: with pixel space loss, full model (MOSAIC) aligns the pixel space color.
Effect of MOSAIC Objective and Multi-Channel Test-Time Optimization. We formulate a naive ablation baseline by extending [16] to scene-level multi-view image generation, averaging each denoising latent prediction $z_{0}$ without our multi-channel inference time optimization for MOSAIC objectives. As shown in Tab. 3, this naive approach yields poor results with the highest KID score and lowest consistency metrics. Fig. 4 reveals significant issues—the second row exhibits severe blurry artifacts (red boxes) and compromised cross-view consistency. In contrast, our full model with pixel space refinement delivers exceptional performance, achieving the lowest KID score, highest consistency metrics, and superior cross-view alignment (green boxes). These results demonstrate that MOSAIC objective optimization is crucial for multi-room environments with significant viewpoint variations, where naive averaging fails to maintain both visual quality and geometric coherence.
Effect of Pixel Space Refinement. Tab. 3 shows how pixel space refinement enhances MOSAIC’s performance. This refinement yields notable improvements in both perceptual quality (lower KID) and cross-view consistency (approaching ground truth metrics). Fig. 4 illustrates this improvement between 3rd and 4th row: texture misalignments (yellow) become properly aligned (green) after refinement.
Effect of Trajectory Key Frame Selection. We evaluate our region-coverage sampling strategy against uniform trajectory sampling in Tab. 3. While image quality and consistency metrics remain comparable between approaches, text-image alignment noticeably decreases with uniform sampling. This occurs because uniform sampling frequently captures frames facing flat walls, challenging VLMs to generate precise captions. Furthermore, pretrained ControlNet models tend to hallucinate textures with such depth inputs, creating semantic misalignments. Our strategic key-frame selection effectively balances information-rich viewpoints with sufficient overlap, yielding more accurate text-conditioned generation without sacrificing consistency.
Table 3: Ablation: Image Quality and Consistency
Image Quality Image Consistency
KID $\downarrow$ CS $\uparrow$ CIQA $\uparrow$ CC $\uparrow$ PSNR $\uparrow$ Ratio $\uparrow$
@@
Figure 5: Second column: Single-view generation creates artifacts (in red) contradicting ground truth depth. Third column: MOSAIC achieves superior depth alignment (in green).
Multi-Views for better RGB-Depth Alignment. Traditional scene-level generation approaches employ warp-and-inpainting strategies that suffer from error accumulation—initial RGB-depth misalignments compound through sequential generation. In contrast, our theoretical analysis in Eq.21 establishes that leveraging multi-view depth conditioning inherently reduces variance, yielding representations that more faithfully reflect ground truth geometry. Tab. 4 empirically validates this theoretical insight: as viewpoint count increases, normalized MSE between ground truth depth and generated image depth (estimated via Depth Anything [38]) demonstrates consistent decline. Fig. 5 visually confirms this phenomenon—single-view conditioning (middle column) produces hallucinated elements that misalign with ground truth geometry (red boxes), while MOSAIC’s two-view approach (right column) maintains precise geometric fidelity. Our analysis further reveals an inflection point in the quality-viewpoint curve, with text-image alignment steadily improving across viewpoints while perceptual quality peaks at 2-3 highly overlapping views. This finding suggests an optimal operating point that balances computational efficiency with generation quality.
@@
NMSE $\downarrow$ CS $\uparrow$ KID $\downarrow$
@@ Acknowledgements
This work was supported in part by NSF IIS-2112633.

### v3

@@ 4.4 Pixel Space Refinement Loss
While our projection loss effectively ensures consistency in latent space, the non-linear nature of the VAE [kingma2022autoencodingvariationalbayes] decoder means that consistency in latent space does not necessarily translate to pixel-space consistency. This discrepancy is particularly pronounced when projection transformations are far from orthogonal. We address this issue through a pixel space refinement process during the final denoising stages. For each view $i$, we decode its predicted latent into pixel space: $\hat{x}^{[i]}=g(\hat{z}_{0}^{[i]})$. It is infeasible to calculate $L_{Proj}$ in pixel space, because back-propagation through multi-channel VAE decoder poses computation difficulties. We propose to warp these decoded images between views and compute representations for each view point by depth-weighted fusion of the overlapped views:
@@
where $\alpha$ controls the selectivity of weighting and $\pi_{ij}$ is the image warping from $j$ view to $i$ view. These optimal pixel-space representations are re-encoded to latent space: ${z}_{0}^{[i]*}=\beta\cdot f({x}^{[i]*})$, where $f$ is the mapping from pixel to latent. The Pixel Space Refinement Loss is defined as:
@@ 4.5 Properties of MOSAIC
5 Experimental Setup
Autonomous Data Collection. To construct our test set, we deployed an indoor-navigation robot that autonomously explored diverse indoor environments. During each trajectory, the robot captured depth maps with its on-board sensors and logged precise camera poses, producing trajectory-aligned inputs compatible with prior work [tang2023mvdiffusion, yang2025scenecraft, hollein2023text2room, controlnetinpaint2023]. We gathered data from 16HM3D [ramakrishnan2021habitatmatterport3ddatasethm3d] and 5 MP3D [chang2017matterport3dlearningrgbddata] scenes, for each scene we collected around 10 independent trajectories. For captioning, we rendered each depth map to a grayscale image and queried the vision–language model GPT-4o [brown2020languagemodelsfewshotlearners] to generate concise, scene-aware descriptions, yielding 2011 (depth, pose, caption) triplets in total.
Baselines. We compare our method with state-of-the-art multi-view generators that leverage geometric priors: MVDiffusion [tang2023mvdiffusion], Warp-Inpaint [controlnetinpaint2023], SceneTex [chen2023scenetexhighqualitytexturesynthesis], and SceneCraft [yang2025scenecraft]. Although our depth maps can be converted to meshes required by SceneTex, running SceneTex itself to synthesize multi-view images for a single scene takes more than 20 hours, making a full scenes evaluation impractical. Therefore, we run SceneTex on an 8-scene subset and report its results separately for fairness. SceneCraft expects semantic layouts with per-object depth; we derive the necessary bounding-box layouts from our dataset. We additionally assess semantic fidelity against Text2Room [hollein2023text2room], a specialist indoor-scene generator.
Evaluation metrics. Our task seeks images that (i) are geometrically consistent across overlapping views, (ii) exhibit high perceptual quality and stylistic coherence, and (iii) faithfully reflect the text prompt. Following MVDiffusion [tang2023mvdiffusion], we measure geometric consistency with Warped PSNR [5596999]—the PSNR after warping one view into another using ground-truth geometry—and its normalized variant against GT, Warped Ratio, where 1.0 indicates perfect alignment. Image quality is assessed with Kernel Inception Distance (KID) [binkowski2018demystifying] and CLIPIQA+ [wang2023exploring]; prompt adherence is quantified by CLIPScore [hessel2021clipscore] and CLIPConsistency [radford2021learning].
Figure 4: Scene-level reconstruction. Fusing the multi-view images generated by MOSAIC with a standard TSDF pipeline produces coherent, textured meshes across diverse indoor rooms.
Table 1: Quantitative evaluation with Kernel Inception Distance (KID), CLIPIQA+, CLIPScore (CS) and CLIPConsistency($C$).
Method KID $\downarrow$ CS $\uparrow$ CIQA $\uparrow$ $C$ $\uparrow$
Warp-Inpaint [controlnetinpaint2023] 0.04646 0.6849 0.6517 29.93
MVDiffusion [tang2023mvdiffusion] 0.03640 0.7016 0.6025 29.41
SceneCraft [yang2025scenecraft] 0.07697 0.7011 0.4531 27.68
Text2Room [hollein2023text2room] 0.08594 0.6618 0.4388 27.34
@@ 5 Experimental Setup
SceneTex [chen2023scenetexhighqualitytexturesynthesis] - small set 0.10536 0.7045 0.6308 28.5740
MOSAIC- small set (ours) 0.06547 0.7141 0.7322 29.5043
@@ 6.1 Evaluation Against Baselines
As illustrated in Fig. 3, MOSAIC generates scenes that are simultaneously more photorealistic and more cross-view–coherent than all competing methods when provided with the same depth inputs. MVDiffusion [tang2023mvdiffusion] exhibits clear view-to-view inconsistencies and is unable to accommodate stylistic prompts such as “in Van Gogh style.” The warp-and-inpainting pipeline [controlnetinpaint2023] produces plausible first frames, yet progressive style drift and error accumulation degrade subsequent views. SceneCraft [yang2025scenecraft], hindered by its NeRF backbone, fails to reconstruct complex layouts and introduces pronounced blur. Text2Room [hollein2023text2room], which also relies on warp-and-inpainting, shows patchwork seams and depth misalignment. SceneTex [chen2023scenetexhighqualitytexturesynthesis] requires over 20 hours to process a single scene, whereas MOSAIC finishes in just a few minutes. Moreover, SceneTex’s score distilling optimization often converges to a single appearance mode, yielding nearly identical textures across seeds; our DDIM sampler preserves diversity under identical prompts. Overall, MOSAIC preserves depth fidelity and stylistic intent across viewpoints and eliminates boundary artifacts through its depth-weighted projection mechanism.
Quantitative Comparisons. Tab. 1 quantitatively compares our approach with Warp-and-Inpainting [controlnetinpaint2023], MVDiffusion [tang2023mvdiffusion], SceneCraft [yang2025scenecraft], Text2Room [hollein2023text2room], and SceneTex [chen2023scenetexhighqualitytexturesynthesis]. MOSAIC achieves the lowest KID ($\downarrow$ 0.0391) and the highest CS ($\uparrow$ 0.7166), CIQA (0.6526), and CC (30.85), outperforming every baseline across all perceptual metrics; it also surpasses SceneTex on the 8-scene subset. Tab. 2 further highlights our superior geometric consistency, recording a Warped PSNR of 25.30 and a Warped Ratio of 0.99—virtually matching ground truth (25.45, 1.00) and eclipsing all alternatives. Although MVDiffusion attains a competitive KID of 0.0364, its Warped PSNR is only 13.58, underscoring its limited multi-view consistency in 3D. In short, MOSAIC delivers state-of-the-art performance in both image fidelity and cross-view alignment.
Figure 5: Qualitative ablation. Columns show matched views for two scenes; rows compare naïve averaging, MOSAIC without $L_{\text{pixel}}$, and full MOSAIC. Boxes indicate identity objects across different viewpoints along generation. Red: blur/ghosts artifacts; orange: inconsistent texture drift; green: full model corrects both.
Table 2: Cross-view geometric consistency analysis.
@@
MVDiffusion [tang2023mvdiffusion] 13.58 0.53
Warp-Inpaint [controlnetinpaint2023] 22.00 0.86
MOSAIC (ours) 25.30 0.99
@@ 6.2 Ablation Study
Effect of MOSAIC Objective and Multi-Channel Test-Time Optimization. We build a naive ablation by extending [kim2024synctweediesgeneralgenerativeframework] to scene-level multi-view image generation, averaging each denoising latent prediction $z_{0}$ without our multi-channel inference time optimization for MOSAIC objectives. As reported in Tab. 3, this variant records the worst KID and the weakest consistency scores. Fig. 5 (second row) further exposes severe blur and cross-view misalignment (red boxes). Conversely, the full MOSAIC model achieves the best KID, the strongest consistency, and visually tight alignment across views (green boxes), confirming that MOSAIC is critical for large viewpoint changes naive averaging collapses both appearance and geometry.
Effect of Pixel Space Refinement. Adding the pixel-space loss $L_{pixel}$ markedly improves performance (Tab. 3), lowering KID and pushing consistency scores close to ground truth. In Fig. 5, boxes track identical objects across different viewpoints: without $L_{pixel}$ (third row) the flower vanishes, the stool disappears, and decorative shapes deform (orange boxes). With $L_{pixel}$ (fourth row) textures remain coherent from different viewpoints (green boxes), demonstrating the benefit of pixel-level supervision.
Effect of Trajectory Key Frame Selection. Tab. 3 contrasts our region-coverage sampling with uniform trajectory sampling. Although overall image quality and consistency are similar, text–image alignment degrades under uniform sampling. Uniformly selected frames often face blank walls, leaving the VLM with insufficient semantic cues; it is also likely to cause pretrained ControlNet [control_net] branches to hallucinate textures. Our key-frame strategy strikes a balance between information-rich viewpoints and sufficient overlap, yielding more accurate text-conditioned results.
Figure 6: RGB–depth alignment Left: reference images at two viewpoints (yellow). Middle: single-view baseline producing geometric artifacts (red). Right: MOSAIC conditioned on multiple depth maps recovers accurate geometry (green).
Multi-Views for better RGB-Depth Alignment. Traditional scene-level generation approaches employ warp-and-inpainting strategies that suffer from error accumulation—initial RGB-depth misalignments compound through sequential generation. By contrast, Eq. 21 shows that conditioning on multiple depth views analytically reduces variance, producing geometry closer to ground truth. To validate this, we select key frames whose viewpoints overlap by more than 50% of their pixels and vary the number $N$ of such views during generation. Tab. 4 corroborates this: the normalized MSE between generated depths (estimated with Depth-Anything [yang2024depthanythingunleashingpower]) and ground-truth depths decreases monotonically as the number of views increases. Fig. 6 visualizes the trend: single-view conditioning (red boxes) hallucinates structures that diverge from ground truth (yellow boxes), whereas multi-view conditioning (green boxes) preserves geometric accuracy. Our analysis further reveals an inflection point in the quality-viewpoint curve, with text-image alignment steadily improving across viewpoints, while perceptual quality peaks at 2-3 highly overlapping views. This finding suggests an optimal operating point that balances computational efficiency with generation quality.
Table 3: Ablation: Image Quality and 3D Consistency
Image Quality Geometry Consistency
KID $\downarrow$ CS $\uparrow$ CIQA $\uparrow$ $C$ $\uparrow$ PSNR $\uparrow$ Ratio $\uparrow$
@@
Views NMSE $\downarrow$ CS $\uparrow$ KID $\downarrow$
@@ Acknowledgements
We are grateful to Yanbo Xu, Zhipeng Bao, Yifan Pu, and Zongtai Li for their helpful comments and discussion. The project was partly supported by NSF IIS-2112633.